Janice has an average of 74 on the first three tests in her algebra class. On the final four tests, she scored 92, 90, 94, and 76. What was her average for the entire year?
Answer by mananth(16949)

Janice has an average of 74 on the first three tests in her algebra class. On the final four tests, she scored 92, 90, 94, and 76. What was her average for the entire year?
The total scores in first three tests = 74*3= 222
she scored 92+90+94+76 = 352 in the next four tests
total score in all tests = 222+352= 574
Total number of test she took for obtaining this score = 7
Average for the whole year = 572/7
=81.71